# Probing the cosmic baryon distribution and the impact of AGN feedback with FRBs in CROCODILE simulation

### v5

@@ 2 Dispersion measure (DM) of FRB
3 LSS Simulations and their Results
3.1 Simulation Setup
@@
Appendices AF provide details on these simulations, contextualizing the role of AGN feedback in our FRB-related results. This supplementary information supports our primary analyses while maintaining a focus on FRB implications in the main text. The following sections present results from LSS simulations, followed by zoom-in cosmological hydrodynamic simulations in Section 6.
@@
3.2 Cosmological Dispersion Measure in LSS Simulations
@@
4 Halo Profile Analysis for DM Contributions of Foreground Halos
@@
4.1 Theoretical Model of Halo Density Profile
@@
4.2 Density profiles derived from LSS simulations
@@
4.3 DM profile as a function of impact parameter
@@
4.4 2D DM Map and 1D Profiles of Individual Halos
@@
5 IGM Baryon Fraction Analysis
In Section 3.2, we constrained an important parameter, $f_{\mathrm{diff}}$, using the DM–$z$ relation for FRBs. This parameter directly reflects the distribution of the "missing baryons" that we are investigating. As the dark matter halos grow, $f_{\mathrm{diff}}$ evolves over time. Additionally, based on numerical simulations, aside from the DM–$z$ relation, we can also estimate the gas mass fraction belonging to the IGM by excluding the gas in the halos, which is often called CGM. Here, we treat the $f_{\mathrm{diff}}$ inferred from the DM–$z$ relation as the observational estimate, and the simulation-derived $f_{\mathrm{IGM}}$ — which excludes the halo (CGM) gas (whose mass dependence is shown in figure 8) — as the theoretical reference or “true value.” In the actual observations, it is usually difficult to subtract the intervening halo contribution, hence $f_{\mathrm{diff}}=f_{\mathrm{IGM}}+f_{\mathrm{halo}}$, as we defined in Sec 2. In some observational papers, the observed $f_{\mathrm{diff}}$ is written as $f_{\mathrm{IGM}}$ assuming that $f_{\mathrm{Halos}}$ is a part of the smooth IGM component. Such an estimate might lead to an overestimate of true $f_{\mathrm{IGM}}$, and a care is needed in its interpretation. In the present work, we attempt to clarify the distinction between $f_{\mathrm{diff}},f_{\mathrm{IGM}}$, and $f_{\mathrm{Halo}}$ in the following sections.
5.1 “Observed” $f_{\mathrm{diff}}$
The calculation of the "observed value" ($f_{\mathrm{diff,obs}}$) follows the same method as in Section 3.2, where we record the constrained $f_{\mathrm{diff,obs}}$ at each redshift when constructing the light cone and connecting the simulation box. We emphasize that the observationally constrained $f_{\mathrm{diff,obs}}$—derived through the integration of DM along light cones—is not strictly equivalent to the instantaneous simulation-based definition of $f_{\mathrm{diff}}(z)$ at a given redshift. The subtle differences are elaborated in Appendix J.
@@ 5 IGM Baryon Fraction Analysis
5.2.1 Using the HMF to Calculate $f_{\mathrm{IGM,true}}$
@@
5.2.2 Directly Subtracting CGM Gas to Calculate $f_{\mathrm{IGM,true}}$
@@
As shown in Figure 9a, all these methods, along with the $f_{\mathrm{IGM}}$ computed via the Press-Schechter HMF approach (Section 5.2.1), converge towards $f_{\mathrm{IGM}}=1$ (gray dashed line) at high redshifts ($z\sim 10$), indicating their consistency in the early Universe when there are few dark matter halos. At lower redshifts, significant differences emerge as halos develop. When removing gas only within $R_{200}$, $f_{\mathrm{IGM}}$ remains the highest and exceeds the Macquart relation constraints, but at $z=1$, it reaches $f_{\mathrm{IGM}}=0.874$ (fiducial) and $0.865$ (NoBH), aligning closely with the Macquart estimate (i.e., using Eq. (8) and regarding $f_{\mathrm{diff}}$ as $f_{\mathrm{IGM}}$). For $R_{\mathrm{cut}}=2R_{200}$, the resulting $f_{\mathrm{IGM}}$ is lower than the Macquart estimate, yielding $f_{\mathrm{IGM}}=0.738$ (fiducial) and $0.733$ (NoBH) at $z=1$. When extending $R_{\mathrm{cut}}$ to $3R_{200}$, the lowest $f_{\mathrm{IGM}}$ values are obtained, with $f_{\mathrm{IGM}}=0.560$ (fiducial) and $0.558$ (NoBH) at $z=1$. These values are obtained only using a single snapshot of CROCODILE simulation at each redshift. The results derived from the FoF halo approach lie between the $R_{200}$ and $2R_{200}$ cases, suggesting that the effective radius defined by the FoF method is within this range. Although the $R_{200}$-based calculation yields $f_{\mathrm{IGM}}$ values closest to the Macquart relation at $z=1$, the overall redshift evolution of $f_{\mathrm{IGM}}$ derived from the FoF halo method exhibits the best agreement with the Macquart relation at low redshifts. This suggests that if CGM is defined within the range of $R_{200}$ to $2R_{200}$, the $f_{\mathrm{diff}}$ constrained by the Macquart relation closely matches the true $f_{\mathrm{IGM}}$ at $z\lesssim 1$.
@@
5.3 Effect of AGN Feedback on Baryon Redistribution
To more precisely analyze the impact of AGN feedback on baryon redistribution, we examine the evolution of halo gas relative to the IGM. Figure 10 shows the redshift evolution of the ratio  [^6] After we define the CGM in Section 5.2.2, the concept of $f_{\mathrm{Halos}}$ can be naturally replaced by $f_{\mathrm{CGM}}$$f_{\mathrm{CGM}}/f_{\mathrm{IGM}}$ for both the fiducial and NoBH models under different $R_{\mathrm{cut}}$ definitions. In general, we see that this ratio is increasing with decreasing redshift, as the halos grow due to gravitational instability. Notably, the observational constraints from L. Connor et al. (2024) are broadly consistent with the predictions adopting $R_{\mathrm{cut}}=R_{200}$, suggesting that this structural definition may best capture the CGM scale probed by current FRB observations. At low redshifts, the ratio is systematically lower in the fiducial case, indicating that AGN feedback expels a portion of halo gas into the IGM. This trend confirms that AGN feedback plays a moderate but non-negligible role in redistributing baryons. This ratio therefore serves as a complementary tracer of AGN-induced baryon redistribution.
@@ 5.3 Effect of AGN Feedback on Baryon Redistribution
5.4 Fitting the Redshift Evolution of $f_{\text{IGM}}$ and $f_{\text{diff}}$
@@
5.5 From FRB Observables to Intrinsic $f_{\mathrm{IGM}}(z)$: A Semi-Analytic Reconstruction Framework
@@
Here, $\langle f_{\mathrm{diff,obs}}(<z)\rangle$ can be inferred from the observation of the Macquart relation up to $z$, which statistically links FRB DMs and redshift. Its analytic form can be approximated using the empirical fitting functions (e.g., C-exp) described in Section 5.4, allowing us to compute the derivative term $\frac{d}{dz}\langle f_{\mathrm{diff,obs}}(<z)\rangle$.
@@
6 Zoom-in Cosmological Hydrodynamic Simulations and their results
@@
6.1 FRBs in Dwarf Galaxies
@@
6.2 FRBs in Spiral Galaxies
@@
For this study, we select a MW-like galaxy with a stellar mass of $1.21\times 10^{11}\,M_{\odot}$ and a star formation rate of $6.6\,M_{\odot}\,\text{yr}^{-1}$, as shown in Table 3, to represent its key characteristics. The analysis method follows the same procedure as in Section 6.1. Results are presented in Figure 13. However, due to the higher mass of the spiral galaxy compared to the dwarf galaxy, the BNS merger timescale is expected to be shorter. Therefore, we adopt a smaller off-center distance $194\,\mathrm{kpc}$.
@@
6.3 FRBs in Galaxy Clusters
@@
For this analysis, we select the largest subhalo near the galaxy cluster halo center as the representative satellite galaxy. This subhalo is located approximately $1.86\,\mathrm{Mpc}$ from the cluster center. The analysis follows the same methodology as that used for spiral galaxies, as outlined in 6.2, but the off-center case is replaced by the satellite case. The results are presented using pink component in Figure 13.
@@ Appendix H Implications of Random Sightline Placement
To assess this potential offset, we refer to our dedicated HH analysis (see Section 6, Figure 13e, f), where the local DM contribution from FRBs varies by environment:
@@
These values estimate the systematic DM contribution omitted by random-LOS sampling. Accordingly, we caution that our DM results represent only the cosmic contribution outside the MW and host galaxies, which are analyzed separately in Section 3.2.
@@ Appendix I Impact of AGN Feedback on 2D DM Distributions in Massive Halos
These results support the findings discussed in Section 4.4 of the main text, emphasizing the varying efficiency of AGN feedback across different halo mass ranges.
@@
In Section 2, we defined the diffuse baryon fraction, $f_{\mathrm{diff}}$, as the sum of those in the IGM and halos. However, both components are redshift-dependent, and thus $f_{\mathrm{diff}}$ evolves with redshift as $f_{\mathrm{diff}}(z)=f_{\mathrm{IGM}}(z)+f_{\mathrm{Halos}}(z)$ (see Figure 10). This relation can be rigorously obtained from cosmological simulations. However, the value of $f_{\mathrm{diff,\,obs}}$ constrained observationally via Equation 8 represents a distinct concept. It reflects the integrated contribution of diffuse gas and halos along the entire sightline up to a given redshift $z$, weighted by the DM contribution of intersected structures.
@@
In the phase-based classification, gas is divided using a temperature threshold of $10^{4}$ K to separate cold and hot components, and a baryon overdensity threshold of $\log_{10}(\rho_{\rm gas}/\bar{\rho}_{b})=3.5$ to distinguish diffuse from dense phases. In contrast, the structure-based classification is based on the CGM subtraction method (Section 5.2.2), in which gas within halos is excluded based on a cutoff radius $R_{\mathrm{cut}}$.

### v6

@@ II Dispersion measure (DM) of FRB
III LSS Simulations and their Results
III.1 Simulation Setup
@@
Appendices AF provide details on these simulations, contextualizing the role of AGN feedback in our FRB-related results. This supplementary information supports our primary analyses while maintaining a focus on FRB implications in the main text. The following sections present results from LSS simulations, followed by zoom-in cosmological hydrodynamic simulations in Section VI.
@@
III.2 Cosmological Dispersion Measure in LSS Simulations
@@
IV Halo Profile Analysis for DM Contributions of Foreground Halos
@@
IV.1 Theoretical Model of Halo Density Profile
@@
IV.2 Density profiles derived from LSS simulations
@@
IV.3 DM profile as a function of impact parameter
@@
IV.4 2D DM Map and 1D Profiles of Individual Halos
@@
V IGM Baryon Fraction Analysis
In Section III.2, we constrained an important parameter, $f_{\mathrm{diff}}$, using the DM–$z$ relation for FRBs. This parameter directly reflects the distribution of the "missing baryons" that we are investigating. As the dark matter halos grow, $f_{\mathrm{diff}}$ evolves over time. Additionally, based on numerical simulations, aside from the DM–$z$ relation, we can also estimate the gas mass fraction belonging to the IGM by excluding the gas in the halos, which is often called CGM. Here, we treat the $f_{\mathrm{diff}}$ inferred from the DM–$z$ relation as the observational estimate, and the simulation-derived $f_{\mathrm{IGM}}$ — which excludes the halo (CGM) gas (whose mass dependence is shown in figure 8) — as the theoretical reference or “true value.” In the actual observations, it is usually difficult to subtract the intervening halo contribution, hence $f_{\mathrm{diff}}=f_{\mathrm{IGM}}+f_{\mathrm{halo}}$, as we defined in Sec II. In some observational papers, the observed $f_{\mathrm{diff}}$ is written as $f_{\mathrm{IGM}}$ assuming that $f_{\mathrm{Halos}}$ is a part of the smooth IGM component. Such an estimate might lead to an overestimate of true $f_{\mathrm{IGM}}$, and a care is needed in its interpretation. In the present work, we attempt to clarify the distinction between $f_{\mathrm{diff}},f_{\mathrm{IGM}}$, and $f_{\mathrm{Halo}}$ in the following sections.
V.1 “Observed” $f_{\mathrm{diff}}$
The calculation of the "observed value" ($f_{\mathrm{diff,obs}}$) follows the same method as in Section III.2, where we record the constrained $f_{\mathrm{diff,obs}}$ at each redshift when constructing the light cone and connecting the simulation box. We emphasize that the observationally constrained $f_{\mathrm{diff,obs}}$—derived through the integration of DM along light cones—is not strictly equivalent to the instantaneous simulation-based definition of $f_{\mathrm{diff}}(z)$ at a given redshift. The subtle differences are elaborated in Appendix J.
@@ V IGM Baryon Fraction Analysis
V.2.1 Using the HMF to Calculate $f_{\mathrm{IGM,true}}$
@@
V.2.2 Directly Subtracting CGM Gas to Calculate $f_{\mathrm{IGM,true}}$
@@
As shown in Figure 9a, all these methods, along with the $f_{\mathrm{IGM}}$ computed via the Press-Schechter HMF approach (Section V.2.1), converge towards $f_{\mathrm{IGM}}=1$ (gray dashed line) at high redshifts ($z\sim 10$), indicating their consistency in the early Universe when there are few dark matter halos. At lower redshifts, significant differences emerge as halos develop. When removing gas only within $R_{200}$, $f_{\mathrm{IGM}}$ remains the highest and exceeds the Macquart relation constraints, but at $z=1$, it reaches $f_{\mathrm{IGM}}=0.874$ (fiducial) and $0.865$ (NoBH), aligning closely with the Macquart estimate (i.e., using Eq. (8) and regarding $f_{\mathrm{diff}}$ as $f_{\mathrm{IGM}}$). For $R_{\mathrm{cut}}=2R_{200}$, the resulting $f_{\mathrm{IGM}}$ is lower than the Macquart estimate, yielding $f_{\mathrm{IGM}}=0.738$ (fiducial) and $0.733$ (NoBH) at $z=1$. When extending $R_{\mathrm{cut}}$ to $3R_{200}$, the lowest $f_{\mathrm{IGM}}$ values are obtained, with $f_{\mathrm{IGM}}=0.560$ (fiducial) and $0.558$ (NoBH) at $z=1$. These values are obtained only using a single snapshot of CROCODILE simulation at each redshift. The results derived from the FoF halo approach lie between the $R_{200}$ and $2R_{200}$ cases, suggesting that the effective radius defined by the FoF method is within this range. Although the $R_{200}$-based calculation yields $f_{\mathrm{IGM}}$ values closest to the Macquart relation at $z=1$, the overall redshift evolution of $f_{\mathrm{IGM}}$ derived from the FoF halo method exhibits the best agreement with the Macquart relation at low redshifts. This suggests that if CGM is defined within the range of $R_{200}$ to $2R_{200}$, the $f_{\mathrm{diff}}$ constrained by the Macquart relation closely matches the true $f_{\mathrm{IGM}}$ at $z\lesssim 1$.
@@
V.3 Effect of AGN Feedback on Baryon Redistribution
To more precisely analyze the impact of AGN feedback on baryon redistribution, we examine the evolution of halo gas relative to the IGM. Figure 10 shows the redshift evolution of the ratio  [^6] After we define the CGM in Section V.2.2, the concept of $f_{\mathrm{Halos}}$ can be naturally replaced by $f_{\mathrm{CGM}}$$f_{\mathrm{CGM}}/f_{\mathrm{IGM}}$ for both the fiducial and NoBH models under different $R_{\mathrm{cut}}$ definitions. In general, we see that this ratio is increasing with decreasing redshift, as the halos grow due to gravitational instability. Notably, the observational constraints from L. Connor et al. (2024) are broadly consistent with the predictions adopting $R_{\mathrm{cut}}=R_{200}$, suggesting that this structural definition may best capture the CGM scale probed by current FRB observations. At low redshifts, the ratio is systematically lower in the fiducial case, indicating that AGN feedback expels a portion of halo gas into the IGM. This trend confirms that AGN feedback plays a moderate but non-negligible role in redistributing baryons. This ratio therefore serves as a complementary tracer of AGN-induced baryon redistribution.
@@ V.3 Effect of AGN Feedback on Baryon Redistribution
V.4 Fitting the Redshift Evolution of $f_{\text{IGM}}$ and $f_{\text{diff}}$
@@
V.5 From FRB Observables to Intrinsic $f_{\mathrm{IGM}}(z)$: A Semi-Analytic Reconstruction Framework
@@
Here, $\langle f_{\mathrm{diff,obs}}(<z)\rangle$ can be inferred from the observation of the Macquart relation up to $z$, which statistically links FRB DMs and redshift. Its analytic form can be approximated using the empirical fitting functions (e.g., C-exp) described in Section V.4, allowing us to compute the derivative term $\frac{d}{dz}\langle f_{\mathrm{diff,obs}}(<z)\rangle$.
@@
VI Zoom-in Cosmological Hydrodynamic Simulations and their results
@@
VI.1 FRBs in Dwarf Galaxies
@@
VI.2 FRBs in Spiral Galaxies
@@
For this study, we select a MW-like galaxy with a stellar mass of $1.21\times 10^{11}\,M_{\odot}$ and a star formation rate of $6.6\,M_{\odot}\,\text{yr}^{-1}$, as shown in Table 3, to represent its key characteristics. The analysis method follows the same procedure as in Section VI.1. Results are presented in Figure 13. However, due to the higher mass of the spiral galaxy compared to the dwarf galaxy, the BNS merger timescale is expected to be shorter. Therefore, we adopt a smaller off-center distance $194\,\mathrm{kpc}$.
@@
VI.3 FRBs in Galaxy Clusters
@@
For this analysis, we select the largest subhalo near the galaxy cluster halo center as the representative satellite galaxy. This subhalo is located approximately $1.86\,\mathrm{Mpc}$ from the cluster center. The analysis follows the same methodology as that used for spiral galaxies, as outlined in VI.2, but the off-center case is replaced by the satellite case. The results are presented using pink component in Figure 13.
@@ Appendix H Implications of Random Sightline Placement
To assess this potential offset, we refer to our dedicated HH analysis (see Section VI, Figure 13e, f), where the local DM contribution from FRBs varies by environment:
@@
These values estimate the systematic DM contribution omitted by random-LOS sampling. Accordingly, we caution that our DM results represent only the cosmic contribution outside the MW and host galaxies, which are analyzed separately in Section III.2.
@@ Appendix I Impact of AGN Feedback on 2D DM Distributions in Massive Halos
These results support the findings discussed in Section IV.4 of the main text, emphasizing the varying efficiency of AGN feedback across different halo mass ranges.
@@
In Section II, we defined the diffuse baryon fraction, $f_{\mathrm{diff}}$, as the sum of those in the IGM and halos. However, both components are redshift-dependent, and thus $f_{\mathrm{diff}}$ evolves with redshift as $f_{\mathrm{diff}}(z)=f_{\mathrm{IGM}}(z)+f_{\mathrm{Halos}}(z)$ (see Figure 10). This relation can be rigorously obtained from cosmological simulations. However, the value of $f_{\mathrm{diff,\,obs}}$ constrained observationally via Equation 8 represents a distinct concept. It reflects the integrated contribution of diffuse gas and halos along the entire sightline up to a given redshift $z$, weighted by the DM contribution of intersected structures.
@@
In the phase-based classification, gas is divided using a temperature threshold of $10^{4}$ K to separate cold and hot components, and a baryon overdensity threshold of $\log_{10}(\rho_{\rm gas}/\bar{\rho}_{b})=3.5$ to distinguish diffuse from dense phases. In contrast, the structure-based classification is based on the CGM subtraction method (Section V.2.2), in which gas within halos is excluded based on a cutoff radius $R_{\mathrm{cut}}$.